OpenStudy (christos):

Solve the inequality x^2+2x>=0

Parth (parthkohli):

\[x(x + 2) \ge 0\]

Parth (parthkohli):

Observe that if you want \(x \times (x + 2)\) nonnegative, both \(x\) and \(x + 2\) are nonnegative.

OpenStudy (christos):

yes

OpenStudy (christos):

hm

OpenStudy (christos):

First time I am solving a 2nd degree inequality :S

OpenStudy (christos):

Could someone help me continue? :)

OpenStudy (anonymous):

I can offer an alternative. Just sketch the graph. That never fails.

OpenStudy (christos):

I need to do it without a graph

OpenStudy (anonymous):

\[x \le -2 \] and \[x \ge0\]

OpenStudy (christos):

How and why?

OpenStudy (anonymous):

Okay, then first find the x-axis intercepts by treating the inequality sign as an equals sign. \[x ^{2}+2x=0\] \[x(x+2)=0\] \[x=-2,0\] Sub in values for \[x<-2\] e.g when x=-3, y=3 Then sub in values for x>0 e.g when x=1, y=3 Since \[x \in (-\infty,-2)\]and \[x \in (0,\infty)\]yield positive answers Therefore \[x ^{2}+2x>=0\] for x=<-2 and x>=0

OpenStudy (mertsj):

Two factors: x and x+2 If the product is positive, x and x+2 must both be positive OR x and x+2 must both be negative. Let us make a picture of the factors and their signs at various points on the number line:

OpenStudy (mertsj):